Deformation and Fracture Mechanics
Answer #1

1.

Ans:

2. - 1 XI 150 =50X1 Up kmoo the ΟΧΙ。 X ( = 구28 (b) Givey 。一 1500 .MPa- SO = 3.5363 X 104 meter

2.

Ans:

The radius of the plastic zone in the direction of the crack can be determined by evaluating the given equation for \theta=0 . r_y=\frac{K_l^2}{2\pi \sigma _y^2}

b) Determine the angle at which the plastic zone is the largest, is a matter of finding a maximum for that equation

dry d (2K2 0 3K2 (2cos --3cos*-) 2K2 (2cos 2(-sino) + 6cos3@sin2)-K2(-sine, 3sinecos2@) -sine 3sinecos)

\frac{dr_y}{d\theta }=\frac{K_l^2}{\pi \sigma _y^2}(-sin\theta+3sin\theta cos^2\frac{\theta}{2} )dr de

1=3 cos^2\frac{\theta}{2}

cos2\sqrt{ \frac{1}{3}}= cos\frac{\theta}{2}\rightarrow \theta=2cos^{-1}(\sqrt{ \frac{1}{3}})

θ = 109.47 deg

3.

Ans:

Sol = 0.4 IC 2. 2 O x IUD